Project

General

Profile

NDOS Commissioning Meeting 2011/06/16

Attendance: Jon, Dennis, Leon, Athans, Jan, Susan, Nick, Dominick, Zeb, Christian, Denis, Peter, Mat, Ken, Enrique

Agenda:

  1. DCS Update & Plans (Athans, Ron, Dennis, ...)
  2. NDOS Status, Hardware Work (Jon, Mat, Denis)
  3. DataCheck update (Susan, Zeb, Nick, Jan, ...)
  4. DAQ/Electronics Update & Plans (Leon, Peter)
  5. AOB (All)

DCS Update & Plans

  • Ron reported to Jon that:
    • ~90% of archiving is done
    • written code to write to the dev. dB, and it seems to work
    • is able to convert raw data to
    • working on set-point and enabling process, hopes to get "something" (eg, not everything, eg, not all channels) working but may need to address the "changing mask" issue (from disabled to enabled)
    • goal is to cool the "golden" APDs on Tuesday
  • Dennis:
    • does not intend to do more in parallel with Ron, but would rather work with Ron
    • this is regardless of the chosen technology, but obviously it would be easier for them if we use ACNet
  • Leon:
    • Athans, Dennis, Ron, Jim K., should discuss what the most optimized decision is.
  • Denis: how many
  • Jon: random

NDOS Status, Hardware Work

  • humidity levels in the past ~2 weeks:
  • Mat:
    • began full detector inspection. All horizontals on 1st diblock is complete, they've removed any APDs with visible damage or gross installation issues.
    • they have observed some corrosion (oxidation) of electronic components on some FEBs. Corrosion found on/near HV connections as well.
  • Leon: it would be good to check connections on PDBs as well.
  • Leon: 12 APDs are in the mail.
  • Mat: ordered more o-rings, glue and adjustable torque wrench.
  • Mat: found some "oil-soaked" APDs. These are strange because there is not a lot of oil visible on the top of the carrier board or on the spacer frame. It's not clear how the oil got there (doesn't look like there was a spill), but Denis mentioned he saw some of these grouped in pairs).

DataCheck Update

DAQ Software/Electronics Update & Plans

  • Peter:
    • looks like the subrun crash problem has been solved (it was actually "fixed" in code, but was not implemented a couple of weeks ago)
    • DCM kernels sometimes go into "panic" when DCMs try to detach/break connections.
  • Kurt:
    • re-prepare run puts RC in a funky/undefined state
    • DL and GT configurations are in the dB; instructions on how to modify these settings are in the DAQ wiki
  • On-mon problem is still present, perhaps getting worse (Enrique).

AOB (All)